# DenseFusion-1M: Merging Vision Experts for Comprehensive Multimodal Perception

### v1

@@ 4.1 Implementation Details
Evaluation Benchmarks. To verify the efficacy of the provided DenseFusion-1M, we adopt these captions during the pre-training stage and follow the setup of LLaVA-1.5 [35] on various visual question answering (VQA) and multi-modality understanding benchmarks for evaluation, such as ScienceQA [39], TextQA [50], VQAv2 [18], GQA [23], SEED [29], MMBench [38], MME [15], POPE [32], MM-Vet [65], that covers a wide range dimensions for evaluating model abilities.
@@ 4.2 Main Results
Experiment Results. (1) Conventionally, Table 2 demonstrates that our meticulous descriptions significantly improve baseline models, providing solid and consistent benefits across all vision-language benchmarks, particularly for text-recognition scenes, e.g. TextVQA. Notably, our dataset originates from the generic LAION, which has no direct connection to the validation domains. Despite this, our strategies outperform ShareGPT4V, which uses images from COCO and VG that share a similar image distribution with the evaluation benchmarks, like VQAv2 and GQA. (2) Additionally, we observe that the potential benefits of our dataset are not fully exploited due to limited input resolutions, making MLLMs challenging to extract hyper-detailed image clues. To address this, we conduct further experiments using the high-resolution MLLM, Scaling on Scales [47] ($\text{S}^{2}$), which performs multi-scale aggregation on high-resolution inputs without increasing the number of image tokens. Even with a fifth of visual tokens of LLaVA-1.6 and requires no additional instruction tuning data, LLaVA-$\text{S}^{2}$ trained by our data achieves better performance than the state-of-the-art LLaVA-1.6 and exhibits higher forward efficiency. Besides, we reproduce LLaVA-$\text{S}^{2}$ using 1.2M pre-training data from ShareGPT4V [8], named ShareGPT4V-$\text{S}^{2}$, and we do not introduce additional supervised fine-tuning data for fair comparisons. Our dataset shows further performance gains compared to the low-resolution version, demonstrating the superiority of our dataset in scenarios requiring hyper-detailed visual elements.
@@ 4.3 Ablation Study
Case Study. We demonstrate caption examples from our perception fusion caption engine and the generalist MLLM LLaVA-1.6 7B [36]. In Fig.5(a), the detected objects help the MLLM focus on individual objects, generating descriptions with more details and attributes. In Fig.5(b), the experts identify tiny bolts on a phone case, OCR with Chinese characters, and LAION captions that provide world knowledge about broader application. This integrated information allows the caption engine to achieve comprehensive image understanding for hyper-detailed captions. Note that even when not all additional information is provided, our caption engine can still focus on producing comprehensive captions, showcasing its robustness. More visualizations are included in the supplementary materials.
@@
Data Efficiency. We conduct the experiment to verify the data efficiency of our high-quality image-text pairs across varying training samples. The experiment results demonstrate our superiority improvements than ShareGPT4V for equivalent data scale. This advantage becomes particularly significant with high-resolution inputs. The experiment indicates that the quality of detailed descriptions and input resolution significantly impact training effectiveness and hyper-detailed captions. As a result, the high-quality image-text data result in a more efficient training manner under the same data scale.
5 Limitation and Discussion
For comprehensive visual perception, we propose perceptual fusion, efficiently integrating the insights from visual experts and constructing the DenseFusion-1M dataset with well-rounded descriptions. Despite its promising results, some issues could be improved: (1) Limited by the current capacity of MLLMs, describing all visible information in an over-complicated image perfectly is still hard to be ensured. (2) The information gathered by visual experts is inevitably noisy; therefore, only high-confidence feedback is retained to ensure accuracy. This can be enhanced by incorporating more sophisticated vision experts. (3) Given the caption engine’s contextual capabilities, only the most crucial visual base models are selected to ensure the full exploitation of contextual information. As the language model’s contextual understanding advances, additional details, such as dense region comprehension, can be progressively integrated.

### v2

@@ 4.1 Implementation Details
Evaluation Benchmarks. To verify the efficacy of the provided DenseFusion-1M, we adopt these captions during the pre-training stage and follow the setup of LLaVA-1.5 [35] on various visual question answering (VQA) and multi-modality understanding benchmarks for evaluation, such as ScienceQA [39], TextQA [50], VQAv2 [18], GQA [23], SEED [29], MMBench [38], MME [15], POPE [32], MM-Vet [65], that covers a wide range dimensions for evaluating model abilities. The metric in Table 2 reflects the individual scores for each benchmark, typically represented as the percentage (%) of correct answers across all questions.
@@ 4.2 Main Results
LLaVA-1.5 LLaMA3-8B 72.3 79.7 63.8 58.7 1553 72.8 69.2 61.8 85.0 34.9
LLaVA-1.5 (Ours) LLaMA3-8B 72.9 80.4 64.4 61.0 1560 73.4 71.6 63.7 85.3 40.0
LLaVA-1.5 Qwen2-7B 72.3 79.8 63.4 57.0 1566 72.9 70.0 62.5 85.7 35.8
LLaVA-1.5 (Ours) Qwen2-7B 73.5 80.5 64.0 58.9 1528 73.5 71.6 63.6 86.0 41.4
@@
Experiment Results. (1) Conventionally, Table 2 demonstrates that our meticulous descriptions significantly improve baseline models, providing solid and consistent benefits across all vision-language benchmarks, particularly for text-recognition scenes, e.g. TextVQA. Notably, our dataset originates from the generic LAION, which has no direct connection to the validation domains. Despite this, our strategies outperform ShareGPT4V, which uses images from COCO and VG that share a similar image distribution with the evaluation benchmarks, like VQAv2 and GQA. (2) Additionally, we observe that the potential benefits of our dataset are not fully exploited due to limited input resolutions, making MLLMs challenging to extract hyper-detailed image clues. To address this, we conduct further experiments using the high-resolution MLLM, Scaling on Scales [47] ($\text{S}^{2}$), which performs multi-scale aggregation on high-resolution inputs without increasing the number of image tokens. Even with a fifth of visual tokens of LLaVA-1.6 and requires no additional instruction tuning data, LLaVA-$\text{S}^{2}$ trained by our data achieves better performance than the state-of-the-art LLaVA-1.6 and exhibits higher forward efficiency. Besides, we reproduce LLaVA-$\text{S}^{2}$ using 1.2M pre-training data from ShareGPT4V [8], named ShareGPT4V-$\text{S}^{2}$, and we do not introduce additional supervised fine-tuning data for fair comparisons. Our dataset shows further gains compared to the low-resolution version, demonstrating our superiority in scenarios requiring hyper-detailed visual elements.
@@ 4.3 Ablation Study
Visual Analysis. We conduct the Visual analysis on specific contribution of visual experts for final description in Figure 6. Besides, We demonstrate caption examples from our perception fusion caption engine and the generalist MLLM LLaVA-1.6 7B [36] in Fig.5. Specially, the detected objects help the MLLM focus on individual objects, generating descriptions with more details and attributes. This integrated information allows the caption engine to achieve comprehensive image understanding for hyper-detailed captions. Note that even when not all additional information is provided, our caption engine can still focus on producing comprehensive captions, showcasing its robustness. More visualizations are included in the supplementary materials.
@@
Figure 6: Visualization on specific contribution of world knowledge, text recognition, and detection model to produce detailed and comprehensive descriptions.
@@
Data Efficiency. We conduct the experiment to verify the data efficiency of our high-quality image-text pairs across varying training samples. The experiment performances (%) demonstrate our superiority improvements than ShareGPT4V for equivalent data scale. This advantage becomes particularly significant with high-resolution inputs. The experiment indicates that the quality of detailed descriptions and input resolution significantly impact training effectiveness and hyper-detailed captions. As a result, the high-quality image-text data result in a more efficient training manner under the same data scale.
@@ 5 Conclusion
6 Acknowledgement
This work was supported by the Program of Beijing Municipal Science and Technology Commission Foundation (No.Z241100003524010), in part by the National Natural Science Foundation of China under Grant 62088102 and the National Key R&D Program of China (2022ZD0116302), in part by AI Joint Lab of Future Urban Infrastructure sponsored by Fuzhou Chengtou New Infrastructure Group and Boyun Vision Co. Ltd, and in part by the PKU-NTU Joint Research Institute (JRI) sponsored by a donation from the Ng Teng Fong Charitable Foundation.